OpenStudy (anonymous):

the limit of 2x-x^2 over 3x+5 as x increases without bound. Anyone could willing to help :(

OpenStudy (anonymous):

minus infinity

OpenStudy (anonymous):

the numerator is a polynomial of degree 2, and the denominator is a polynomial of degree 1, so as x gets large this looks like \[\frac{-2x^2}{x}=-x\]

OpenStudy (anonymous):

actually i should have written \[\frac{-x^2}{3x}=-\frac{x}{3}\] but of course as x goes to infinity the whole thing goes to minus infinity

OpenStudy (anonymous):

thnaks satellite

OpenStudy (anonymous):

yw